OpenStudy (anonymous):

Suppose you have a simple circuit that includes a resistance device of 25 ohms and the current flowing through it is 5 amps. a) Solve for the voltage necessary showing all your steps. b) Solve for the power showing all your steps.

OpenStudy (anonymous):

please!

OpenStudy (anonymous):

Power = (current squared)(resistance) Power = I^2 R = (5)^2 (25) = (25)(25)= 625 watts